Straight-Line Depreciation Rates Convert each of the following estimates of useful life to a straight-line depreciation rate, stated as a percentage:

(a) 10 years, (b) 8 years, (c) 25 years, (d) 40 years, (e) 5 years, (f) 4 years, (g) 20 years. If required, round your answers to one decimal place. Years Percentage a. 10 years % b. 8 years % c. 25 years % d. 40 years % e. 5 years % f. 4 years % g. 20 years % PreviousNext
Answer #1
Straight-Line Depreciatiion rate =1 / no. of useful life in years
(a)For 10 Years-1 / 10 Years =10%
(b)For 8 Years-1 / 8 Years =12.5%
©For 25 Years-1 / 25 Years =4%
(d)For 40 Years-1/40 Years =2.5%
(e)For 5 Years-1/5 Years =20%
(f)For 4 Years-1/4 Years =25%
(g)For 20 Years-1/20 Years =5%